Solve for x algebraically: 5^(2x)- 4(5)^x=-3
Answer by stanbon(75887)

5^(2x)- 4(5)^x=-3
----------------------
(5^x)^2 -4(5^x) +3 = 0
Let 5^x = w
--------------
w^2-4w+3 = 0
(w-3)(w-1) = 0
w = 3 or w = 1
--------------------
Replace w by 5^x:
5^x = 3 or 5^x = 1
x = log(3)/log(5) or x = 0
x = 0.682606... or x = 0
======================
Cheers,
Stan H.